# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

### v2

@@ Abstract.
The Beilinson–Bloch conjecture is a generalization of the Birch and Swinnerton-Dyer conjecture, which relates the ranks of Chow groups of smooth projective varieties over global fields to the order of vanishing of $L$-functions. We prove the conjecture for certain classes of non-isotrivial varieties over ${\mathbb{F}}_{q}(t)$, including some cubic threefolds. We deduce the Birch and Swinnerton-Dyer conjecture for their intermediate Jacobians, and use it to establish new cases of the Tate conjecture over finite fields. We also prove some additional results on the arithmetic of these intermediate Jacobians.
@@ 1. Introduction
Let $X$ be a smooth projective variety over a global field $k$. The Beilinson–Bloch conjecture ([Bloch1984], [Beilinson1987]) predicts that the rank of the homologically trivial part of the codimension-$i$ Chow group of $X$ is equal to the order of vanishing at $s=i$ of the Hasse–Weil $L$-function $L(H^{2i-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s)$. For $X$ an abelian variety and $i=1$, this is equivalent to the (weak) Birch and Swinnerton-Dyer conjecture for $X$. In general, few cases of the Beilinson–Bloch conjecture have been established unconditionally, particularly over number fields, where the $L$-function is not even known in general to be well-defined at $s=i$. We recall some existing results on the conjecture in subsection 4.2.
@@
In this paper we will be concerned with the Beilinson–Bloch conjecture in positive characteristic. Tate [Tate1965] and Artin–Tate [Tate1966b] recognized in the 1960s that the Birch and Swinnerton-Dyer conjecture over global function fields was closely related to the Tate conjecture for divisors on varieties over finite fields. They conjectured a special value formula for the zeta function of a smooth projective surface over a finite field, and predicted that under mild hypotheses, when the surface was fibered over a curve, the validity of that formula was equivalent to the strong Birch and Swinnerton–Dyer conjecture for the Jacobian of the generic fiber. This equivalence, termed Artin–Tate’s conjecture (d), was proved in the early 2000s by combining results of Artin–Tate, Grothendieck [Grothendieck1968], Milne [Milne1986], and Kato–Trihan [KT2003]. See [LRS2022] for some additional historical background, and two alternative proofs of conjecture (d).
@@
Geisser in [Geisser2021, Theorem 1.1] established a higher-dimensional generalization of conjecture (d), which applies (under mild hypotheses) to any smooth proper variety $X$ over a finite field which is fibered over a curve. Geisser’s theorem can be reformulated as saying that the Tate conjecture for divisors on $X$ is equivalent to the Tate conjecture for divisors on the generic fiber, together with the Birch and Swinnerton-Dyer conjecture for the Albanese variety of the generic fiber. In [Broe2025, Corollary 5.6], we generalized part of Geisser’s result to cycles of higher codimension, giving a sufficient condition for the Beilinson–Bloch conjecture to hold for a variety over a function field. From this we deduced some new cases of the conjecture for isotrivial varieties (similar isotrivial examples were known previously, such as those given in [Jannsen2007, Remark 5.2] and [Kahn2021, Corollary 1]). Here a variety over a global function field $K/{\mathbb{F}}_{q}$ is called isotrivial if its base change to $\overline{K}$ is defined over $\overline{{\mathbb{F}}}_{q}$.
@@
It remains to show that this method can prove cases of the Beilinson–Bloch conjecture in higher codimension for interesting non-isotrivial varieties. Since our criterion for the conjecture requires some control over the bad reduction of a variety over a function field, it is natural to seek examples where the singularities of the special fibers are as mild as possible. We are thus led to consider Lefschetz pencils. Recall that a Lefschetz pencil for a smooth projective variety $X$ is roughly a generically smooth family of hyperplane sections of $X$, parametrized by $\mathbb{P}^{1}$, such that each singular member of the family (over the algebraic closure) is smooth away from one ordinary double point singularity. They have long played a structural role in algebraic and arithmetic geometry, notably in Deligne’s papers on the Weil conjectures ([Deligne1974], [Deligne1980]).
@@
This establishes the conjecture for a class of smooth cubic threefolds over ${\mathbb{F}}_{q}(t)$. Under some stronger hypotheses on $X$, we use results of Jannsen [Jannsen2007] to compute all motivic cohomology groups (with rational coefficients) of the generic fiber of $f$, in terms of certain simpler invariants related to étale cohomology (theorem 6.12). We also prove cases of the Beilinson–Bloch conjecture for cubic hypersurfaces of higher dimension (theorem 6.5), and establish the finiteness of the Albanese kernel for a class of surfaces over ${\mathbb{F}}_{q}(t)$ (corollary 6.9). The varieties covered by these results are shown to be non-isotrivial in many cases (remark 6.11): for instance, the cubic threefolds of theorem 1.1 are never isotrivial.
@@
The proof of theorem 1.1 relies on the Tate conjecture for cubic fourfolds over finite fields, established unconditionally in [Charles2013], [Madapusi2015], and [AKPW2025]. Note that a similar Lefschetz pencil argument was used by Charles–Pirutka [CP2015] in their proof of the integral Tate conjecture for smooth cubic fourfolds in characteristic $\neq 2,3$.
@@
Smooth cubic threefolds over $\mathbb{C}$ have long been studied through their intermediate Jacobians, which are principally polarized abelian fivefolds constructed from the Hodge structures of the threefolds. Clemens–Griffiths [CG1972] famously took this approach to show that smooth complex cubic threefolds are never rational, a theorem soon extended by Murre [Murre1973] to base fields of characteristic not 2. Later work of Murre [Murre1985], or rather its generalization by Achter–Casalaina-Martin–Vial [ACMV2023], permits a definition of the intermediate Jacobian of a smooth cubic threefold $X$ over an arbitrary field. This abelian variety satisfies a certain universal property, and admits a concrete geometric interpretation: it is isomorphic to the Albanese variety of the Fano surface of lines on $X$ [Ciurca2024, Proposition 5.54]. Ciurca in loc. cit. has also announced a proof of the irrationality of smooth cubic threefolds in characteristic 2: his manuscript is currently only available as a preprint, and we do not rely on it for any of our results here.
@@
There is a certain relation between the Chow motive of a smooth cubic threefold over any field and that of its intermediate Jacobian (proposition 3.4). It formally follows that the Beilinson–Bloch conjecture for a smooth cubic threefold over a global field is equivalent to the Birch and Swinnerton-Dyer conjecture for its intermediate Jacobian. From this and [KT2003], we deduce the strong form of BSD for the intermediate Jacobians of the cubic threefolds treated in theorem 1.1. Since the rational $\ell$-adic Tate modules of these abelian varieties are controlled by the well-understood monodromy of Lefschetz pencils, we can make several other observations about their arithmetic. In particular, they have everywhere semistable reduction, and their geometric isogeny classes are simple and non-isotrivial (proposition 7.1). We moreover explicitly compute their $L$-functions, and give an example of Mordell–Weil rank 28. In future work, we aim to undertake a more focused study of their BSD invariants.
@@
We conclude the paper by proving some new cases of the Tate conjecture over finite fields (remark 7.4). Specifically, we consider a smooth cubic threefold $X$ over a global function field of characteristic $\geq 3$, and assume that $X$ satisfies the Beilinson–Bloch conjecture (which holds unconditionally for the cubic threefolds of theorem 1.1). Using the Tate conjecture for the Fano surface $F$ of $X$, proved by Roulleau [Roulleau2013], and the Birch and Swinnerton-Dyer conjecture for the intermediate Jacobian of $X$, combined with the backwards direction of the aforementioned theorem of Geisser, we establish the Tate conjecture for a certain auxiliary threefold over the constant field, obtained by spreading out $F$. It is interesting that here the shortest path between two cases of the Tate conjecture over finite fields seems to pass through the domain of function fields, and we hope that the argument demonstrates some utility of the general Beilinson–Bloch conjecture in positive characteristic.
@@
Section 2 recalls some results on the structure of Chow groups of certain varieties, particularly cubic hypersurfaces. Section 3 introduces tools from the theory of Chow motives, and uses them to relate the motive of a smooth cubic threefold with that of its intermediate Jacobian. In section 4 we state Jannsen’s generalized Tate conjecture and the Beilinson–Bloch conjecture, and give sufficient conditions for them. Section 5 collects results from SGA 7 [SGA7-II] and Weil II [Deligne1980] on the monodromy of Lefschetz pencils. These permit an explicit computation (under mild hypotheses) of the $L$-function associated with the middle cohomology of the generic fiber of a Lefschetz pencil over ${\mathbb{F}}_{q}$, in terms of the zeta function of the total space of the pencil. Section 6 proves cases of the Beilinson–Bloch conjecture for generic fibers of certain Lefschetz pencils, essentially via geometric arguments. Section 7 deduces BSD for the intermediate Jacobian of a smooth cubic threefold over ${\mathbb{F}}_{q}(t)$ which arises from a Lefschetz pencil of a cubic fourfold, and uses it to obtain cases of the Tate conjecture over ${\mathbb{F}}_{q}$. Here we also employ the earlier results on monodromy of Lefschetz pencils to describe the arithmetic of the intermediate Jacobian, and compute its $L$-function.
@@
The letter $k$ will denote a field (arbitrary unless stated otherwise), and $\overline{k}$ will denote a choice of algebraic closure. We let $G_{k}$ denote the absolute Galois group of $k$. A variety over $k$ is a finite-type separated integral $k$-scheme. By $\ell$ we will always mean a fixed prime not equal to the characteristic of the base field. Our conventions for motives follow André’s book [Andre2004], and also match those of the Stacks project [stacks-project, Tag 0FG9]. By default, all Chow groups and motives have rational coefficients, and are taken with respect to rational equivalence. For a smooth projective variety $X$, the Chow group modulo homological equivalence (with respect to $\ell$-adic cohomology) is denoted by $\operatorname{CH}^{i}(X)_{\hom}$, and the homologically trivial subgroup of $\operatorname{CH}^{i}(X)$ is denoted by $\operatorname{CH}^{i}(X)_{0}$. By $\operatorname{CH}_{i}(X)$ we mean the Chow group graded by dimension of cycles, rather than codimension. If $R$ is a commutative ring, the Chow group with coefficients in $R$ is denoted by $\operatorname{CH}^{i}(X)_{R}$. The category of motives over $k$ is denoted by $\mathcal{M}(k)$. The tensor unit in this category is denoted by $\mathbf{1}$. The motive of a smooth projective scheme $Y$ is denoted by $\mathfrak{h}(Y)$. The dual of a motive $M$ is denoted by $M^{\vee}$, and the dual of a morphism $f:M\to N$ of motives is denoted by $f^{\vee}:N^{\vee}\to M^{\vee}$. The $n$-fold Tate twist of $f$ is denoted by $f(n):M(n)\to N(n)$. Our sign convention for Tate twists of motives is such that $\mathfrak{h}(\mathbb{P}^{1})\cong\mathbf{1}\oplus\mathbf{1}(-1)$. Subscripts on schemes, motives, or morphisms thereof denote base change. By $H^{i}(Y_{\overline{k}},{\mathbb{Q}}_{\ell})$ we mean $\ell$-adic étale cohomology. For a finite field ${\mathbb{F}}_{q}$, $\operatorname{Frob}_{q}$ denotes the geometric Frobenius in $G_{{\mathbb{F}}_{q}}$, which sends $x\mapsto x^{1/q}$.
@@ Acknowledgments
I thank Jeff Achter, Niranjan Ramachandran, Padma Srinivasan and Doug Ulmer for helpful conversations.
@@ 2. Preliminaries on algebraic cycles
We begin by collecting some known results on the Chow groups of certain varieties. Our focus is on cubic hypersurfaces (whose Chow groups are also surveyed in [Huybrechts2023, Section 7.7.4]). Throughout the paper we will freely use the following standard lemma.
@@
For any finite-type scheme $X$ over $k$ and any algebraic field extension $L/k$, the pullback map $\operatorname{CH}_{i}(X)\to\operatorname{CH}_{i}(X_{L})$ is injective. When $X$ is additionally smooth, the pullback is also injective when $L/k$ is an arbitrary extension.
@@
The former claim is [stacks-project, Tag 0FH8]. The latter follows from the argument of [LS2025, Lemma 2.1]. This uses the refined Gysin homomorphism of [Fulton1998, Chapter 6.2]; since a closed immersion of finite-type smooth schemes is a regular embedding [Fulton1998, Appendix B.7.2], the Gysin homomorphism is well-defined in the generality needed here. ∎
@@
In particular, the proposition applies when $X$ is a smooth cubic hypersurface of dimension $\geq 2$ and $\neq 3$: then $X_{\Omega}$ is unirational [Kollar2002, Theorem 1], hence rationally chain connected [Fulton1998, Example 10.1.6], so $\operatorname{CH}_{0}(X_{\Omega})_{\mathbb{Z}}\cong{\mathbb{Z}}$.
@@
If $k$ is an algebraic extension of a finite field, then for any smooth cubic hypersurface $X$ over $k$ of dimension $\geq 3$, $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$.
@@
It suffices to assume that $k=\overline{k}$. Then by [Shen2019, Theorem 1.7], there exists a surjective map $\operatorname{CH}_{0}(F(X))\to\operatorname{CH}_{1}(X)$. Here $F(X)$ is the Fano variety of lines on $X$, which is smooth and connected by [AK1977, Theorems 1.10 and 1.17]. Proposition 2.4 implies that $\operatorname{CH}_{0}(F(X))\cong{\mathbb{Q}}$, whence the result. ∎
We note in passing that [Shen2019, Theorem 1.7] has been generalized to cycles of higher dimension in [Lyu2025, Theorem 0.1], though we will not use this.
@@
Proposition 2.6 ([ELV1997, Theorem 4.6]).
Let $Z$ be a closed subscheme of $\mathbb{P}^{n}_{k}$, defined by $r$ homogeneous equations of respective degrees $d_{1}\geq...\geq d_{r}\geq 2$. If $j\in{\mathbb{Z}}$ and either $d_{1}\geq 3$ or $r\geq j+1$, then whenever
@@
we have $\operatorname{CH}_{j^{\prime}}(Z)\cong{\mathbb{Q}}$ for $0\leq j^{\prime}\leq j$.
For example, in the case of a cubic hypersurface $Z$ of dimension $\geq 5$ over any field, we find that $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$.
@@
Let $k$ be a finitely generated field and $X$ a smooth cubic fourfold over $k$. If $\operatorname{char}(k)\neq 2$, or $k$ is finite of characteristic 2, then the Tate conjecture holds for $X$.
@@
See [Charles2013, Corollary 6], [Madapusi2015, Theorem 5.14], and [AKPW2025, Theorem 4.9]. ∎
Remark 2.8.
We believe the Tate conjecture for smooth cubic fourfolds over finitely generated fields of characteristic 2 should follow from the case of a finite base field, via standard specialization arguments. However, this is beyond our present scope.
When $X$ is a smooth cubic hypersurface of dimension $d$, its middle Betti number $b_{d}(X)$ is computed in [Huybrechts2023, Corollary 1.1.12]: when $d=2,$ $3,$ or $4$, $b_{d}(X)$ is respectively 7, 10, or 23. We will freely use these values throughout the paper without further comment. We also note the classical result that a smooth cubic surface over an algebraically closed field is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points [Huybrechts2023, Proposition 4.2.4].

### v3

@@ Abstract.
The Beilinson–Bloch conjecture is a generalization of the Birch and Swinnerton-Dyer conjecture, which relates the ranks of Chow groups of smooth projective varieties over global fields to the order of vanishing of $L$-functions. We prove the conjecture for certain classes of non-isotrivial varieties over ${\mathbb{F}}_{q}(t)$, including some cubic threefolds and fivefolds. We deduce the Birch and Swinnerton-Dyer conjecture for their intermediate Jacobians, and use it to establish new cases of the Tate conjecture over finite fields. We also prove further results on the arithmetic of these intermediate Jacobians. To that end, we show that a few classes of varieties over an arbitrary field have motive of abelian type, generalizing previously known examples over the complex numbers.
@@ 1. Introduction
Let $X$ be a smooth projective variety over a global field $k$. The Beilinson–Bloch conjecture ([9], [8]) predicts that
(1.1) $\dim_{{\mathbb{Q}}}\operatorname{CH}^{i}(X)_{0}=\operatorname*{ord}\limits_{s=i}L(H^{2i-1}(X_{\overline{k}},{\mathbb{Q}}_{\ell}),s).$
[…]
@@
In this paper we will be concerned with the Beilinson–Bloch conjecture in positive characteristic. Tate [91] and Artin–Tate [93] recognized in the 1960s that the Birch and Swinnerton-Dyer conjecture over global function fields was closely related to the Tate conjecture for divisors on varieties over finite fields. They conjectured a special value formula for the zeta function of a smooth projective surface over a finite field, and predicted that under mild hypotheses, when the surface was fibered over a curve, the validity of that formula was equivalent to the strong Birch and Swinnerton-Dyer conjecture for the Jacobian of the generic fiber. This equivalence, termed Artin–Tate’s Conjecture (d), was proved in the early 2000s by combining results of Artin–Tate, Artin–Grothendieck [41], Milne [76], and Kato–Trihan [57]. See [68] for some additional historical background, and two alternative proofs of Conjecture (d).
@@
Geisser in [39, Theorem 1.1] established a higher-dimensional generalization of Conjecture (d), which applies (under mild hypotheses) to any smooth proper variety $X$ over a finite field which is fibered over a curve. Geisser’s theorem can be reformulated as saying that the Tate conjecture for divisors on $X$ is equivalent to the Tate conjecture for divisors on the generic fiber, together with the Birch and Swinnerton-Dyer conjecture for the Albanese variety of the generic fiber. Building on work of Jannsen [47], in [14, Corollary 4.15] we generalized part of Geisser’s result to cycles of higher codimension, giving a sufficient condition for the Beilinson–Bloch conjecture to hold for a variety over a function field. From this we deduced some new cases of the conjecture for isotrivial varieties (similar isotrivial examples were known previously, such as those given in [50, Remark 13.6.2] and [51, Corollary 1]). Here a variety over a global function field $K/{\mathbb{F}}_{q}$ is called isotrivial if its base change to $\overline{K}$ is defined over $\overline{{\mathbb{F}}}_{q}$.
@@
It remains to show that this method can prove cases of the Beilinson–Bloch conjecture in higher codimension for interesting non-isotrivial varieties. Since our criterion for the conjecture requires some control over the bad reduction of a variety over a function field, it is natural to seek examples where the singularities of the special fibers are as mild as possible. We are thus led to consider Lefschetz pencils. Recall that a Lefschetz pencil for a smooth projective variety $X$ is roughly a generically smooth family of hyperplane sections of $X$, parametrized by $\mathbb{P}^{1}$, such that each singular member of the family (over the algebraic closure) is smooth away from one ordinary double point singularity. They have long played a structural role in algebraic and arithmetic geometry, notably in Deligne’s papers on the Weil conjectures ([28], [29]).
@@
This establishes the conjecture for a class of smooth cubic threefolds over ${\mathbb{F}}_{q}(t)$. Under some stronger hypotheses on $X$, we use results of Jannsen [50] to compute all motivic cohomology groups (with rational coefficients) of the generic fiber of $f$, in terms of certain simpler invariants related to étale cohomology (theorem 6.12). We also prove cases of the Beilinson–Bloch conjecture for cubic hypersurfaces of higher dimension (theorem 6.5), and establish the finiteness of the Albanese kernel for a class of surfaces over ${\mathbb{F}}_{q}(t)$ (corollary 6.9). The varieties covered by these results are shown to be non-isotrivial in many cases (remark 6.11): for instance, the cubic threefolds of theorem 1.1 are never isotrivial.
@@
The proof of theorem 1.1 relies on the Tate conjecture for cubic fourfolds over finite fields, established unconditionally in [16], [72], and [7]. Note that a similar Lefschetz pencil argument was used by Charles–Pirutka [15] in their proof of the integral Tate conjecture for smooth cubic fourfolds in characteristic $\neq 2,3$.
@@
Smooth cubic threefolds over $\mathbb{C}$ have long been studied through their intermediate Jacobians, which are principally polarized abelian fivefolds constructed from the Hodge structures of the threefolds. Clemens–Griffiths [19] famously took this approach to show that smooth complex cubic threefolds are never rational, a theorem soon extended by Murre [80] to base fields of characteristic not two. Later work of Murre [81], as generalized by Achter–Casalaina-Martin–Vial [3], permits a definition of the intermediate Jacobian of a smooth cubic threefold $X$ over an arbitrary field. This abelian variety satisfies a certain universal property, and admits a concrete geometric interpretation: it is isomorphic to the Albanese variety of the Fano surface of lines on $X$ [18, Proposition 5.54]. Ciurca in loc. cit. has also announced a proof of the irrationality of smooth cubic threefolds in characteristic two: his manuscript is currently only available as a preprint, and we do not rely on any of its results in the present work.
@@
There is a certain relation between the Chow motive of a smooth cubic threefold over any field and that of its intermediate Jacobian (proposition 3.12). It formally follows that the Beilinson–Bloch conjecture for a smooth cubic threefold over a global field is equivalent to the Birch and Swinnerton-Dyer conjecture for its intermediate Jacobian. From this and [57], we deduce the strong form of BSD for the intermediate Jacobians of the cubic threefolds treated in theorem 1.1. Since the rational $\ell$-adic Tate modules of these abelian varieties are controlled by the well-understood monodromy of Lefschetz pencils, we can make several other observations about their arithmetic. In particular, they have everywhere semistable reduction, and their geometric isogeny classes are simple and non-isotrivial (proposition 7.1). We moreover explicitly compute their $L$-functions, and give an example of Mordell–Weil rank 28. In future work, we aim to undertake a more focused study of their BSD invariants.
We also show that a few other classes of smooth complete intersections over any field have associated abelian varieties, well-defined up to isogeny, which serve as motivic analogues of intermediate Jacobians (corollary 3.21). In characteristic zero, this follows from combining results of Vial [99] and Achter–Casalaina-Martin–Vial [2]: the proof over a general base field then proceeds via a standard specialization argument. Similarly to the case of cubic threefolds, the Beilinson–Bloch conjecture for a smooth cubic fivefold is equivalent to BSD for the associated “intermediate Jacobian”. From theorem 6.5, we thus deduce BSD for the intermediate Jacobians of certain cubic fivefolds over ${\mathbb{F}}_{q}(t)$, and give examples of Mordell–Weil rank $\geq 86$.
@@
We conclude the paper by proving some new cases of the Tate conjecture over finite fields (theorem 7.5). Specifically, we consider a smooth cubic threefold $X$ over a global function field, and assume that $X$ satisfies the Beilinson–Bloch conjecture (which holds unconditionally for the cubic threefolds of theorem 1.1). Using the Tate conjecture for the Fano surface $F$ of $X$, proved in [86] and [24, Remark 4.2], and the Birch and Swinnerton-Dyer conjecture for the intermediate Jacobian of $X$, combined with the backwards direction of the aforementioned theorem of Geisser, we establish the Tate conjecture for a certain auxiliary threefold over the constant field, obtained by spreading out $F$.
@@
Section 2 recalls some results on the structure of Chow groups of certain varieties, particularly cubic hypersurfaces. Section 3 introduces tools from the theory of Chow motives, and uses them to relate the motive of a smooth cubic threefold with that of its intermediate Jacobian. We prove a similar relation for smooth cubic fivefolds, and a few other complete intersections. The latter proof uses some standard specialization techniques for Chow motives, which nevertheless have further applications which we did not find in the literature. Namely, we use them to generalize to arbitrary base fields several classes of varieties which are known over $\mathbb{C}$ to have finite-dimensional motive. These include the Fano varieties of lines on smooth cubic threefolds and fivefolds.
In section 4 we state Jannsen’s generalized Tate conjecture and the Beilinson–Bloch conjecture, and give sufficient conditions for them. Section 5 collects results from SGA 7 [27] and Weil II [29] on the monodromy of Lefschetz pencils. These permit an explicit computation (under mild hypotheses) of the $L$-function associated with the middle cohomology of the generic fiber of a Lefschetz pencil over ${\mathbb{F}}_{q}$, in terms of the zeta function of the total space of the pencil. Other than the definitions, the content of this section finds its primary application in section 7.
Section 6 proves cases of the Beilinson–Bloch conjecture for generic fibers of certain Lefschetz pencils, essentially via geometric arguments. These are enabled by the simple geometry of the singular fibers of Lefschetz pencils of cubic hypersurfaces (lemma 6.3). Section 7 deduces BSD for the intermediate Jacobian of a smooth cubic threefold over ${\mathbb{F}}_{q}(t)$ which arises from a Lefschetz pencil of a cubic fourfold, and uses it to obtain cases of the Tate conjecture over ${\mathbb{F}}_{q}$. Here we also employ the earlier results on monodromy of Lefschetz pencils to describe the arithmetic of the intermediate Jacobian, and compute its $L$-function. We prove analogous statements for the intermediate Jacobians of certain smooth cubic fivefolds.
@@
The letter $k$ will denote a field (arbitrary unless stated otherwise), and $\overline{k}$ will denote a choice of algebraic closure. We let $G_{k}=\operatorname{Aut}(\overline{k}/k)$ denote the absolute Galois group of $k$. A variety over $k$ is a finite-type separated integral $k$-scheme. Unless stated otherwise, $\ell$ will always mean a fixed prime not equal to the characteristic of the base field. Our conventions for motives follow André’s book [5], and also match those of the Stacks project [95, Tag 0FG9]. By default, all Chow groups and motives have rational coefficients, and are taken with respect to rational equivalence. For a smooth projective variety $X$, the Chow group modulo homological equivalence (with respect to $\ell$-adic cohomology) is denoted by $\operatorname{CH}^{i}(X)_{\hom}$, and the homologically trivial subgroup of $\operatorname{CH}^{i}(X)$ is denoted by $\operatorname{CH}^{i}(X)_{0}$. By $\operatorname{CH}_{i}(X)$ we mean the Chow group graded by dimension of cycles, rather than codimension. When $i<0$, both $\operatorname{CH}^{i}(X)$ and $\operatorname{CH}_{i}(X)$ are defined to be zero. The category of motives over $k$ is denoted by $\mathcal{M}(k)$. The tensor unit in this category is denoted by $\mathbf{1}$. The motive of a smooth projective scheme $Y$ is denoted by $\mathfrak{h}(Y)$. The dual of a motive $M$ is denoted by $M^{\vee}$, and the dual of a morphism $f:M\to N$ of motives is denoted by $f^{\vee}:N^{\vee}\to M^{\vee}$. The $n$-fold Tate twist of $f$ is denoted by $f(n):M(n)\to N(n)$. Our sign convention for Tate twists of motives is such that $\mathfrak{h}(\mathbb{P}^{1})\cong\mathbf{1}\oplus\mathbf{1}(-1)$. Subscripts on schemes, motives, or morphisms thereof denote base change. We will sometimes consider motives and Chow groups modulo other adequate equivalence relations, and also denote this by a subscript (e.g. if $M$ is a motive, then $M_{\operatorname{num}}$ denotes $M$ considered modulo numerical equivalence). If $R$ is a commutative ring, the Chow group with coefficients in $R$ is denoted by $\operatorname{CH}^{i}(X)_{R}$, and the category of motives with coefficients in $R$ is denoted by $\mathcal{M}(k)_{R}$. By $H^{i}(Y_{\overline{k}},{\mathbb{Q}}_{\ell})$ we mean $\ell$-adic étale cohomology. For a finite field ${\mathbb{F}}_{q}$, $\operatorname{Frob}_{q}$ denotes the geometric Frobenius in $G_{{\mathbb{F}}_{q}}$, which sends $x\mapsto x^{1/q}$.
We caution that some of our references (e.g. [99] and [100]) adopt a covariant convention for motives, meaning that they work in the opposite of our category $\mathcal{M}(k)$, and use the opposite sign convention for Tate twists. We will translate to our preferred conventions when using their results. See [52, Section 14.1.2] for a guide to the covariant convention.
@@ Acknowledgments
I thank Jeff Achter, Niranjan Ramachandran, Shubhankar Sahai, Padma Srinivasan, Doug Ulmer, and Charles Vial for helpful conversations.
@@ 2. Preliminaries on algebraic cycles
We begin by collecting some known results on the Chow groups of certain varieties. Our focus is on cubic hypersurfaces (whose Chow groups are also surveyed in [46, Section 7.7.4]). Throughout the paper we will freely use the following standard lemma.
@@
For any finite-type scheme $X$ over $k$ and any algebraic field extension $L/k$, the pullback map $\operatorname{CH}_{i}(X)\to\operatorname{CH}_{i}(X_{L})$ is injective. When $X$ is additionally smooth and separated, the pullback is also injective when $L/k$ is an arbitrary extension.
@@
This follows from [95, Tag 0FH8] and [14, Lemma 2.2]. ∎
@@
In particular, the proposition applies when $X$ is a smooth cubic hypersurface of dimension $\geq 2$ and $\neq 3$: then $X_{\Omega}$ is unirational [61, Theorem 1], hence rationally chain connected [38, Example 10.1.6], so $\operatorname{CH}_{0}(X_{\Omega})_{\mathbb{Z}}\cong{\mathbb{Z}}$.
The Fano variety of lines encodes much of the geometry of a cubic hypersurface. We note some of its basic properties.
Proposition 2.4 ([4, Theorems 1.10 and 1.16]).
For $X$ a smooth cubic hypersurface over $k$ of dimension $\geq 3$, the Fano variety of lines on $X$ is smooth, projective, and geometrically connected.
@@
If $k$ is an algebraic extension of a finite field, then for any smooth cubic hypersurface $X$ over $k$ of dimension $\geq 3$, we have $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$.
@@
It suffices to assume that $k=\overline{k}$. Then by [89, Theorem 1.7], there exists a surjective map $\operatorname{CH}_{0}(F(X))\to\operatorname{CH}_{1}(X)$, where $F(X)$ is the Fano variety of lines on $X$. By our assumption on $k$, the group of $k$-points of $\operatorname{Alb}(F(X))$ is torsion, so proposition 2.5 implies that $\operatorname{CH}_{0}(F(X))\cong{\mathbb{Q}}$. As any nonzero effective class in $\operatorname{CH}_{1}(X)$ is nontrivial, we must have $\operatorname{CH}_{1}(X)\cong{\mathbb{Q}}$. ∎
We note in passing that [89, Theorem 1.7] has been generalized to cycles of higher dimension in [70, Theorem 0.1], though we will not use this.
@@
Proposition 2.7 ([33, Theorem 4.6]).
Let $Z$ be a closed subscheme of $\mathbb{P}^{n}_{k}$, defined by $r$ homogeneous equations of respective degrees $d_{1}\geq...\geq d_{r}\geq 2$. Fix $j\geq 0$. If either $d_{1}\geq 3$ or $r\geq j+1$, then whenever
@@
we have $\operatorname{CH}_{j^{\prime}}(Z)\cong{\mathbb{Q}}$ for $0\leq j^{\prime}\leq j$. If $d_{1}=2$ and $1\leq r\leq j$, then whenever
$r(j+2)\leq n+r-j-1,$
the same conclusion holds.
For example, in the case of a cubic hypersurface $Z$ of dimension $\geq 5$ over any field, we find that $\operatorname{CH}_{1}(Z)\cong{\mathbb{Q}}$.
@@
Let $k$ be a finitely generated field and $X$ a smooth cubic fourfold over $k$. If $\operatorname{char}(k)\neq 2$, or $k$ is finite of characteristic two, then the Tate conjecture holds for $X$.
@@
See [16, Corollary 6], [72, Theorem 5.14], and [7, Theorem 4.9]. ∎
When $X$ is a smooth cubic hypersurface of dimension $d$, its middle Betti number $b_{d}(X)$ is computed in [46, Corollary 1.1.12]: when $d=2,$ $3,$ 44, 55, or $6$, $b_{d}(X)$ is respectively 7, 10, 23, 42, or 87. We will freely use these values throughout the paper without further comment. We also note the classical result that a smooth cubic surface over an algebraically closed field is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points [46, Proposition 4.2.4].